The expected real rate of interest is 0.6%, actual inflation over the last year was 3%, and expected inflation over the next year is 7.4%. What is the current level of nominal interest rates (in %) predicted by the Fisher equation? Round to 0.01%. E.g., if your answer is 3.145%, record it as 3.15.
Answer #1

As per the Fisher equation

(1+r) = (1+R)/ (1+h) where r = real rate

R= nominal rate

h= Inflation

(1+0.006)= (1+R)/ (1+0.03)

(1+R) = 1.006*1.03

=1.0362

R= Nominal rate = 3.62%